Operations Report March 8th

SOL: 15

Name of person filing report: Mathéo Fouchet

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

ROVERS

Spirit rover used: No

Hours: 157.7

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Opportunity rover used: Yes

Hours: 86.3

Beginning charge: 100%

Ending charge: 77%

Currently charging: Yes

Curiosity rover used: Yes

Hours: 170.3

Beginning charge: 100%

Ending charge: 76%

Currently charging: Yes

Perseverance rover used: No

Hours: 228.6

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

General notes and comments: Nothing to report

Summary of Hab operations:

WATER USE: 15 gallons

Water (static tank): 262 gallons

Water (loft tank): 40 gallons

Water Meter: 01563385 units

Static to Loft Pump used: Yes

Static tank pipe heater: On

Static tank heater: On

Toilet tank emptied: No

Summary of internet: Nothing to report

Summary of suits and radios: There is a crack under one of the screws present on the red ring of a suit’s helmet. We are doing our best to take care of the material that has been provided to us. I may need help to fix it, I thought about using glue to re-stick both parts together. Let me know if you have any advice on how to fix this issue.

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations:

Dual split: Off

SOC: 92% at 6:32pm

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report